# Position: Foundation Agents as the Paradigm Shift for Decision Making

### v1

@@ C.2 Implementation Details
Pretraining and Fine-tuning. We jointly pretrain our agent on 5 DMC tasks and 5 Atari games using the AdamW optimizer (Loshchilov & Hutter, 2017) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1e-8$. For pretraining, the learning rate is originally set to be 1e-5 with linear warm-up and cosine schedule decay. We use batch size of 512 and weight decay of 0.1 for pretraining. For fine-tuning, the learning rate keeps constant of 1e-4 with batch size 256 and dropout rate 0.1. During fine-tuning, we model and predict returns and rewards in addition to actions, and perform expert action inference as introduced in (Lee et al., 2022). We use context length $L=50$ for both pretraining and fine-tuning [^2] Our code is based on https://github.com/microsoft/smart.. The size of our pretrained model and corresponding computational requirements can be referred to Table 2. Since this is simply a case study of potential foundation agent in applications of robotics and game play, we only provide the visualization of fine-tuning performance in Figure 2 to demonstrate its feasibility and effectiveness in decision making scenarios. We do not report other experimental results and compare with baselines as done in original research. In addition, the case study only involves the first two steps in the roadmap of foundation agents as specified in Figure 1. We consider aligning the pretrained model in this case study with LLMs in future work.

### v2

@@ C.2 Implementation Details
Pretraining and Fine-tuning. We jointly pretrain our agent on 5 DMC tasks and 5 Atari games using the pretraining objective of next action prediction $-logP_{\theta}(a_{t}|\tau_{0:t-1},o_{t})$ via standard mean-squared error loss for continuous actions or cross-entropy loss for discrete actions. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1e-8$. For pretraining, the learning rate is originally set to be 1e-5 with linear warm-up and cosine schedule decay. We use batch size of 512 and weight decay of 0.1 for pretraining. For fine-tuning, the learning rate keeps constant of 1e-4 with batch size 256 and dropout rate 0.1. During fine-tuning, we model and predict returns and rewards in addition to actions using the learning objective $-logP_{\phi}(R_{t},a_{t},r_{t}|\tau_{0:t-1},o_{t})$, and perform expert action inference as introduced in (Lee et al., 2022). We use context length $L=50$ for both pretraining and fine-tuning [^2] Our code is based on https://github.com/microsoft/smart.. The size of our pretrained model and corresponding computational requirements can be referred to Table 2. Since this is simply a case study of potential foundation agent in applications of robotics and game play, we only provide the visualization of fine-tuning performance in Figure 2 to demonstrate its feasibility and effectiveness in decision making scenarios. We do not report other experimental results and compare with baselines as done in original research. In addition, the case study only involves the first two steps in the roadmap of foundation agents as specified in Figure 1. We consider aligning the pretrained model in this case study with LLMs in future work.